OpenStudy (anonymous):

for some reason i am having trouble remembering how to integrate from a negative bound to positive. my example is the integral from -2 to 2 of 4-x^2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

now since 1/ is an even function u can reduce the limits from 0 to 2 and re write it as, 2(4-x^2)... now continue normally.....can u?

OpenStudy (anonymous):

Take the integral of the equation \[4x-1/3x^3\] from -2 to 2 put in the higher bound and subtract the lower bound 16/3-(-16/3) =32/3 True that the equation is an even function, but only with even functions do you need to split the bounds

OpenStudy (anonymous):

I got 64/3... 8 -8/3 - (-8 -8/3)=...

OpenStudy (anonymous):

or (if integrate from 0 - 2) 2*(8-8/3) = 64/3